LOW TEMPERATURE MILLIMETER WAVE DOUBLE RESONANCE OF METHYL FLUORIDE

\underline{C. D. BALL}, M. M. BEAKY, AND F. C. DE LUCIA, Department of Physics, The Ohio State University, 174 W. 18th Avenue, Columbus, OH 43210.

We have recently undertaken an investigation of molecular collision processes employing time resolved double resonance techniques. A millimeter wave - millimeter wave double resonance experiment becomes possible at low temperatures where, unlike at room temperature, a millimeter wave pump can induce large deviations from the thermal population distribution. We use the collisional cooling method to cool molecules to very low temperatures (around 4 K). Resonances, which result from quasi-bound states, dominate low temperature pressure broadening cross sections and also contribute to low temperature rotational energy transfer. In our first experiments, we have measured state-to-state energy transfer rates for methyl fluoride (collisionally cooled by helium). We compare the pressure dependence of these rates with measured pressure broadening data.